### v2

# FlexTaF: Enhancing Table Reasoning with Flexible Tabular Formats

Abstract
The table reasoning task aims to answer the question according to the given table. Currently, using Large Language Models (LLMs) is the predominant method for table reasoning. Most existing methods employ a fixed tabular format to represent the table, which could limit the performance. Given that each instance requires different capabilities and models possess varying abilities, we assert that different instances and models suit different tabular formats. We prove the aforementioned claim through quantitative analysis of experimental results, where different instances and models achieve different performances using various tabular formats. Building on this discussion, we propose FlexTaF-Single and FlexTaF-Vote to enhance table reasoning performance by employing flexible tabular formats. Specifically, (i) FlexTaF-Single trains a classifier to predict the most suitable tabular format based on the instance and the LLM. (ii) FlexTaF-Vote integrates the results across different formats. Our experiments on WikiTableQuestions and TabFact reveal significant improvements, with average gains of $2.3\%$ and $4.8\%$ compared to the best performance achieved using a fixed tabular format with greedy decoding and self-consistency decoding, thereby validating the effectiveness of our methods [^1] Our code and prompts are available at https://github.com/zhxlia/FLEXTAF..
1 Introduction
Table reasoning is a crucial task in natural language processing that aims to automatically extract and infer information from tables (Dong et al., 2022). In this task, a model needs to answer questions based on the information contained in the table, with each question-table pair referred to as an instance. Given the superior commonsense and logical reasoning capabilities of Large Language Models (LLMs), researchers increasingly utilize them for table reasoning, which has become the mainstream method (Chen, 2023; Zhang et al., 2024d; Dong & Wang, 2024). Therefore, we focus on how to solve the table reasoning task with LLMs in this paper.
Figure 1: The table reasoning performance varies with different tabular formats. The List format is convenient for sequential indexing, while the Database format facilitates the search for columns that meet specific conditions.
Some previous works enhance the table reasoning performance by designing prompts (Cheng et al., 2023; Ye et al., 2023; Zhang et al., 2024e; Lee et al., 2024), such as Chain-of-Table Wang et al. (2024), which solves table reasoning tasks by prompting the LLM to iteratively reason with natural language and programs. While the above methods provide a fixed tabular format in their prompts, Sui et al. (2024); Singha et al. (2023); Deng et al. (2024) argue that different table reasoning tasks have different most suitable tabular formats. However, existing works primarily analyze from the task aspect without considering that solving different instances demands distinct reasoning skills Shi et al. (2020); Chen et al. (2023); Cheng et al. (2023); Liu et al. (2024a); Chen et al. (2024b). For example, as shown in Figure 1, solving instance (a) involves reasoning with sequential indexing, which is facilitated by the List format. In contrast, solving instance (b) requires identifying columns that satisfy specific conditions, making the Database format more suitable. Therefore, using a fixed tabular format for all instances could limit performance. Additionally, models vary in their reasoning capabilities (Cao et al., 2023; Bhandari et al., 2024; Zhang et al., 2024a), therefore the most suitable tabular format could differ for each model.
Based on the above discussion, we focus on the impact of tabular formats on the table reasoning performance of LLMs from the following two aspects. (i) We claim that different instances and different LLMs require distinct most suitable tabular formats [^2] We refer to the suitable tabular format as the one that enables the model to correctly solve a given instance.. (ii) We propose to enhance the table reasoning performance by predicting the most suitable format or assembling the results from multiple tabular formats.
First of all, we discuss that distinct tabular formats are suitable for different instances and LLMs. We conduct exploratory experiments utilizing different tabular formats and LLMs on table reasoning datasets. The experimental results present that table reasoning performance varies significantly with different formats and LLMs. Based on the above analysis, we propose FlexTaF, which includes FlexTaF-Single and FlexTaF-Vote, to enhance table reasoning performance through flexible tabular formats. FlexTaF-Single identifies the most suitable tabular format based on the instance and the LLM by training the classifier. FlexTaF-Vote determines the final answer by voting on results obtained from multiple formats. In comparison, although FlexTaF-Single requires training data but only infers once, FlexTaF-Vote is training-free but with expensive inference costs.
To demonstrate the effectiveness of our methods, we conduct experiments on two mainstream table reasoning datasets: WikiTableQuestions (Pasupat & Liang, 2015) and TabFact (Chen et al., 2020). Compared to the best performances achieved by using a fixed format with greedy decoding and self-consistency decoding Wang et al. (2023), FlexTaF-Single and FlexTaF-Vote show average improvements of $2.3\%$ and $4.8\%$ respectively with comparable inference costs, confirming the effectiveness of our methods. Further analysis experiments reveal that some instances can only be correctly solved using a specific format, proving that the most suitable tabular formats for different instances are distinct.
Our contributions are as follows:
  1. 1.
    We claim that the most suitable tabular formats for different instances and LLMs are distinct.
  2. 2.
    We propose FlexTaF, which includes FlexTaF-Single and FlexTaF-Vote, to enhance table reasoning performance by utilizing flexible tabular formats.
  3. 3.
    Our experimental results indicate that FlexTaF-Single and FlexTaF-Vote achieve average improvements of $2.3\%$ and $4.8\%$ respectively, over the best results obtained using fixed formats with greedy decoding and self-consistency decoding at comparable inference costs, demonstrating the effectiveness of our methods.
2 Discussion on Impact of Tabular Formats
We first claim that the most suitable tabular formats for different instances and LLMs are distinct. Since resolving different instances requires diverse abilities Chen et al. (2023); Cheng et al. (2023); Chen et al. (2024b), it is necessary to tailor tabular formats for each instance accordingly. Additionally, the capabilities of different models vary (Cao et al., 2023; Bhandari et al., 2024; Zhang et al., 2024a), resulting in different tabular formats suitable for different LLMs. We further discuss from the perspective of experimental results in this section.
We select five popular formats, including Markdown, Dict, List, Pandas, and Database formats, following previous works Singha et al. (2023); Cheng et al. (2023); Wang et al. (2024); Liu et al. (2024a). Detailed descriptions of these formats are provided in the supplementary material. Subsequently, we quantitatively analyze how tabular format affects the table reasoning performance of LLMs from the aspects of the instance and the model respectively.
2.1 The Relationship between Instance and Tabular Format
We first claim that different instances suit different tabular formats. Table 4 presents the percentage of instances that can only be correctly resolved by each tabular format. It can be observed that, with the same fixed LLM, each tabular format is uniquely suited to certain instances, with some formats correctly solving over $20\%$ of the instances exclusively. The results indicate that, for a given LLM, the most suitable tabular formats vary according to the specific instances.
Figure 2: The overview of FlexTaF. FlexTaF-Single consists of two steps: (i) Classification: A classifier we trained predicts the most suitable tabular format based on the given instance and model. (ii) Reasoning: Using the predicted format, the LLM solves the instance by representing the table accordingly. FlexTaF-Vote consists of two steps: (i) Reasoning: Various formats are employed to represent the table and facilitate reasoning with the LLM, resulting in multiple answers. (ii) Vote: The final answer is determined using a voting mechanism.
2.2 The Relationship between Model and Tabular Format
In this subsection, we discuss that different models suit different tabular formats. Table 2 demonstrates that the most suitable tabular format varies across models and the performance gap arises due to different formats. For instance, Llama3 Meta (2024) exhibits significantly better performance with the Markdown format compared to other formats, while DeepSeek-Coder Guo et al. (2024) performs better with Dict and Database formats than with Markdown. We employ a Chi-square test to demonstrate significant differences in the distribution of the most suitable tabular formats across different models, as detailed in the supplementary material.
3 Methodology
In this section, we introduce FlexTaF, which consists of FlexTaF-Single and FlexTaF-Vote. The overview of FlexTaF is shown in Figure 2.
3.1 Task Definition
FlexTaF focuses on the table reasoning task, which can be formally defined as follows: Given an instance $I$ comprising a natural language question $Q$ and a table $T$, the table reasoning task aims to derive the corresponding answer $\hat{A}=\texttt{M}(\texttt{F}(T),Q)$, where M represents the model and F denotes the tabular format. To effectively solve the table reasoning task, the probability that the generated answer $\hat{A}$ matches the gold answer $A^{*}$ should be maximized.
3.2 FlexTaF-Single
As discussed in §2, it is essential to determine the most suitable tabular format, so FlexTaF-Single achieves this by training a classifier to identify the most suitable format.
3.2.1 Classification
Classification aims to predict the most suitable tabular format from a set of candidate formats based on the instance and the model. It can be formally expressed as $\hat{\texttt{F}}=\texttt{CLS}_{\texttt{M}}(I),\hat{\texttt{F}}\in\texttt{F}_{\texttt{M},I}^{*}=\{\texttt{F}|\texttt{M}(\texttt{F}(T),Q)=A^{*}\}$.
Training Data Collection
To train the classifier to predict the most suitable tabular format for the given instance $I$ and the LLM M, we annotate the training data with M. Specifically, for each instance in the training set, we utilize all candidate tabular formats to reason respectively and evaluate the correctness of each answer. We then collect the set of most suitable tabular formats for each instance in the training data, denoted as $\{\texttt{F}_{\texttt{M},I_{t}}^{*}\}$, for which M can correctly reason with the format on the instance $I_{t}$, and use these as the training data for our classifier. Additionally, we take a data filtering strategy to remove instances from the training set where more than half of the candidate formats are correct or where none are correct, as such instances do not effectively highlight differences between the tabular formats.
Learning Objective
Since there could be multiple formats in $\texttt{F}_{\texttt{M},I}^{*}$, we apply a multi-label classification training method (Godbole & Sarawagi, 2004; Tsoumakas & Katakis, 2007; Herrera et al., 2016), where each label denotes a tabular format. Moreover, we utilize a binary relevance method (Godbole & Sarawagi, 2004) for multi-label classification. Specifically, the tabular formats are serialized and transformed into binary vectors for each instance. During training, we adopt Binary Cross-Entropy Loss, normalized by the number of instances $N$, as follows:
$\texttt{L}(\hat{y},y)=-\frac{1}{N}\sum_{i=1}^{N}\sum_{r=1}^{|\texttt{F}|}[y_{ir}\texttt{log}(\hat{y}_{ir})+(1-y_{ir})\texttt{log}(1-\hat{y}_{ir})]$ (1)
[…]
Predicting Tabular Format
After obtaining the classifier, we predict the most suitable tabular format $\hat{\texttt{F}}$. We regularize the predicted scores of all formats and select the format with the highest probability as our classification result $\hat{\texttt{F}}$.
3.2.2 Reasoning
After predicting the most suitable table format $\hat{\texttt{F}}$, we utilize the predicted format for table reasoning. Specifically, we represent the table with the predicted format $\hat{\texttt{F}}$ in the prompt and employ the LLM to derive the final answer.
3.3 FlexTaF-Vote
FlexTaF-Single requires manual data annotation, which could be difficult to obtain, so we propose FlexTaF-Vote to obtain the final answer by integrating the results from multiple tabular formats, inspired by Wang et al. (2023); Qin et al. (2023); Luo et al. (2024).
3.3.1 Reasoning
We first construct multiple table reasoning prompts, representing the table with different formats in each prompt. Then we employ these prompts to reason with the LLM, obtaining multiple answers accordingly.
3.3.2 Vote
We retain the most consistent result across multiple results by adopting a voting mechanism, which can be formally expressed as the following equation.
$\hat{A}=\texttt{argmax}_{A}\sum_{r=1}^{|\texttt{F}|}\mathds{1}(A_{r}=A)$ (2)
[…]
3.4 Comparison
To better employ our two methods, we examine their application scenarios. (iFlexTaF-Single is ideal for scenarios where high-efficiency online reasoning is required, although it necessitates prior training data annotation and classifier training. (iiFlexTaF-Vote suits scenarios where annotated data is unavailable while maintaining a certain tolerance for real-time reasoning efficiency.
Tabular Format WikiTQ TabFact
Llama3 DeepSeek-Coder Llama3 DeepSeek-Coder
8B 70B 6.7B 33B 8B 70B 6.7B 33B
Markdown $47.7$ $63.3$ $32.2$ $31.2$ $75.2$ $86.4$ $60.4$ $63.6$
Dict $43.0$ $56.4$ $25.6$ $53.6$ $65.5$ $80.0$ $63.9$ $78.0$
List $30.5$ $56.3$ $19.2$ $50.8$ $57.4$ $77.5$ $63.7$ $75.4$
Pandas $39.2$ $52.3$ $39.7$ $48.8$ $47.8$ $73.6$ $62.5$ $75.9$
Database $31.0$ $48.2$ $41.8$ $45.5$ $65.0$ $75.0$ $70.7$ $76.3$
FlexTaF-Single $50.5$ $69.1$ $46.3$ $54.5$ $77.0$ $87.1$ $70.9$ $78.3$
$\Delta$ $+2.8$ $+5.8$ $+4.5$ $+0.9$ $+1.5$ $+0.7$ $+2.2$ $+0.3$
Table 1: The accuracy of reasoning using a fixed tabular format with greedy decoding and FlexTaF-Single, across four LLMs on WikiTableQuestions (WikiTQ) and TabFact. The best performance for each LLM and dataset is marked in bold. $\Delta$ denotes the improvement of FlexTaF-Single relative to the best performance of using a fixed format with greedy decoding for each LLM and dataset.
4 Experiments
4.1 Settings
4.1.1 Datasets
We use WikiTableQuestions (Pasupat & Liang, 2015) and TabFact (Chen et al., 2020) datasets to evaluate FlexTaF. WikiTableQuestions is a mainstream dataset for the table Question Answering task, containing diverse questions across various domains, which requires answering the question based on the table. TabFact is a prominent dataset for the table fact verification task, which needs to determine whether a claim is entailed or refuted by the table.
Tabular Format WikiTQ TabFact
Llama3 DeepSeek-Coder Llama3 DeepSeek-Coder
8B 70B 6.7B 33B 8B 70B 6.7B 33B
Markdown $49.1$ $62.8$ $32.1$ $29.7$ $75.2$ $86.7$ $60.9$ $63.6$
Dict $44.8$ $58.1$ $28.7$ $59.7$ $67.9$ $80.9$ $71.9$ $82.4$
List $35.4$ $59.7$ $27.5$ $55.6$ $59.8$ $78.0$ $66.9$ $78.0$
Pandas $41.4$ $56.2$ $43.9$ $54.4$ $56.0$ $74.7$ $67.2$ $79.2$
Database $35.2$ $48.6$ $42.5$ $46.4$ $69.5$ $76.0$ $70.7$ $77.2$
FlexTaF-Vote $55.7$ $69.9$ $51.4$ $60.9$ $80.3$ $88.5$ $77.9$ $84.4$
$\Delta$ $+6.6$ $+7.1$ $+7.5$ $+1.2$ $+5.1$ $+1.8$ $+7.2$ $+2.0$
Table 2: The accuracy of reasoning using a fixed tabular format with self-consistency decoding Wang et al. (2023) and FlexTaF-Vote, across four LLMs on WikiTQ and TabFact. The best performance for each LLM and dataset is marked in bold. $\Delta$ denotes the improvement of FlexTaF-Vote relative to the best performance of using a fixed format with self-consistency for each LLM and dataset.
4.1.2 Metric
We employ accuracy as the evaluation metric for WikiTableQuestions and TabFact, following the previous works (Pasupat & Liang, 2015; Chen et al., 2020), and use accuracy to evaluate the performance of classification. Accuracy measures the model ability to generate the correct answer, which is achieved only when the predicted result exactly matches the correct answer. Since FlexTaF-Single aims to identify the most suitable format, we adopt accuracy to assess whether the top format predicted by the classifier is in the suitable formats of the instance.
4.1.3 Models
For reasoning, we employ Llama3 (Meta, 2024) and DeepSeek-Coder (Guo et al., 2024), and for classification, we use ELECTRA-Large Clark et al. (2020). Llama3 and DeepSeek-Coder are popular open-source LLMs for their outstanding performance across various tasks. We choose ELECTRA-Large due to its superior performance in language comprehension and question-answering tasks compared to other pre-trained models of similar size (Clark et al., 2020).
4.1.4 Implementation Details
For the WikiTableQuestions and TabFact datasets, we adopt Markdown, Dict, List, Pandas, and Database as tabular formats, which are commonly used in previous works (Singha et al., 2023; Cheng et al., 2023; Ye et al., 2023; Wang et al., 2024) and generally have high performance across datasets with different models. To enhance table reasoning performance, we utilize the Chain-of-Thought (CoT) prompt (Wei et al., 2022) for Markdown and the Program-of-Thought (PoT) prompt (Chen et al., 2023; Gao et al., 2023) for other formats. In addition, we apply 4-shot prompts and 2-shot prompts respectively, since the questions in WikiTableQuestions are more challenging. To improve the performance on WikiTableQuestions, we use different demonstrations for each tabular format, and we explore the performance of FlexTaF with unified demonstrations in §4.3.1. Detailed prompts are provided in the supplementary material. We train a tabular format classifier for each LLM on each dataset. We provide detailed training information in the supplementary material. For sampling in self-consistency decoding Wang et al. (2023), we set $temperature=0.1$ which can bring optimal performance across most formats within $temperature\leq 0.8$. To ensure a fair comparison with FlexTaF-Vote, we set $sampling\_n=5$.
4.2 Main Experiment
Table 1 and Table 2 compare FlexTaF-Single with using a fixed format with greedy decoding, and FlexTaF-Vote with using a fixed format with self-consistency Wang et al. (2023), respectively. We observe that: (iFlexTaF-Single and FlexTaF-Vote surpass the best results achieved by the fixed format with greedy decoding and self-consistency, by an average of $2.3\%$ and $4.8\%$ respectively, with comparable computational costs. (ii) Compared to flexible tabular formats, the fixed format restricts table reasoning performance even with self-consistency. Moreover, self-consistency with Markdown improves slightly or even decreases compared to greedy decoding. This is attributed to self-consistency instability in CoT prompting Chen et al. (2024a); Renze & Guven (2024), and diminished instruction following ability at higher sampling temperatures Zeng et al. (2024); Peeperkorn et al. (2024). Additionally, the tables reveal that:
The improvement of FlexTaF on the more challenging dataset is more significant.
Both FlexTaF-Single and FlexTaF-Vote achieve performance improvement across datasets, underscoring their efficacy. Specifically, the performance on WikiTQ is significantly better compared to the simpler TabFact. Despite higher classification accuracy on TabFact (Table 5), FlexTaF-Single shows limited improvement due to the already high-performance baseline. Moreover, FlexTaF-Vote exhibits less improvement on TabFact due to the greater overlap between simper instances correctly solved by different formats (see supplementary material).
FlexTaF-Single shows superior performance on the general model compared to the code model.
FlexTaF-Single and FlexTaF-Vote demonstrate efficacy across diverse models. Notably, FlexTaF-Single exhibits greater improvement on the general models than on the code models. The tabular formats suitable for the code model are closely related to code, resulting in smaller differences and consequently limited classifier performance.
FlexTaF-Vote consistently surpasses FlexTaF-Single.
Although the performance of FlexTaF-Single is constrained by classification accuracy (see Table 5), it demonstrates high reasoning efficiency which only infers once. In contrast, FlexTaF-Vote achieves superior table reasoning performance across various LLMs and datasets, because instances could be resolved correctly by multiple formats, and errors produced by different tabular formats exhibit diversity. However, FlexTaF-Vote is less efficient in reasoning.
4.3 Analysis
We select Llama3-8B for subsequent analytical experiments due to its high reasoning efficiency. Moreover, we conduct most experiments on WikiTQ, because it encompasses more diverse questions (Pasupat & Liang, 2015; Chen et al., 2020; Dong et al., 2022; Shi et al., 2020).
4.3.1 Is the impact of tabular formats due to the different demonstrations?
MD Dict List PD DB $\text{Acc}_{\text{max}}-\text{Acc}_{\text{min}}$
$47.7$ $43.0$ $30.5$ $39.2$ $16.1$ $31.6$
Table 3: The accuracy on the WikiTQ dataset using Llama3-8B with greedy decoding, employing unified demonstrations in the prompts, which are different from the prompts used in the main experiments. MD denotes Markdown, PD denotes Pandas, and DB denotes Database.
We conduct experiments with unified demonstrations in the prompts for each format and present the results in Table 3. Specifically, we adopt the demonstrations of the same question across different tabular formats, manually annotating the rationale or program accordingly. Detailed prompts are present in the supplementary material. Table 3 indicates that the tabular formats continue to significantly affect performance, with the performance gap widening when using unified demonstrations.
4.3.2 Are the tabular formats suitable for different instances different?
Figure 3: The overlap between instances solved by tabular formats achieved by Llama3-8B on WikiTQ. The values represent the proportion of instances that can be solved by the tabular format corresponding to the vertical axis, within the instances solvable by the format on the horizontal axis.
Model Scale MD Dict List PD DB
Llama3 8B $24.6$ $7.1$ $4.6$ $5.9$ $8.3$
70B $16.0$ $1.7$ $2.2$ $1.8$ $4.3$
DeepSeek-Coder 6.7B $24.7$ $6.7$ $3.7$ $11.0$ $13.6$
33B $15.6$ $3.9$ $3.3$ $4.4$ $7.5$
Table 4: The percentage of instances that can only be correctly solved by one tabular format, in all instances that the tabular format can correctly solve.
We analyze the overlap between instances correctly solved by each format and the proportion of instances that can only be solved by a specific format, as shown in Figure 3 and Table 4. Figure 3 illustrates that: (i) The overlaps between instances correctly solved by each format are all $\leq 80\%$, indicating the distinctions among formats. (ii) The Dict and Pandas formats exhibit higher overlap due to their superior performance (see Table 2), while the DB and List formats have lower overlap. The overlaps between instances solved by different formats using four LLMs on WikiTQ and TabFact are provided in the supplementary material.
Table 4 presents that: (i) Certain instances can only be accurately solved using a specific format, highlighting the differences between formats. (ii) Markdown and Database formats resolve a greater proportion of instances compared to Dict, List, and Pandas formats due to their distinct structures compared to the programming formats. Figure 3 and Table 4 claim that different instances suit different tabular formats.
4.3.3 How does the number of candidate tabular formats affect FlexTaF?
MD+ Dict+ PD+ DB+ List4545454550505050555555556060606047.747.747.747.750.150.150.150.150.450.450.450.450.450.450.450.450.550.550.550.547.747.747.747.748.548.548.548.553.953.953.953.956.256.256.256.255.755.755.755.7accuracyFlexTaF-SingleFlexTaF-Vote
Figure 4: The accuracy of FlexTaF-Single and FlexTaF-Vote using Llama3-8B on WikiTQ with different numbers of candidate tabular formats, as additional candidate tabular formats are added from left to right.
We perform experiments using varying numbers of candidate formats and present results in Figure 4, where formats are added in descending order of performance. We observe the following: (i) The performance of FlexTaF-Single gradually stabilizes as the number of candidate formats increases. The classification performance does not always improve due to the increased difficulty with a higher number of labels Wang et al. (2021); Audibert et al. (2024). (iiFlexTaF-Vote varies greatly with the increase in the number of formats due to its reliance on the performance of each format. In particular, FlexTaF-Vote does not exceed FlexTaF-Single when Markdown and Dict are candidates, as it selects from two different answers only by comparing probabilities, which do not accurately indicate correctness Wang et al. (2023); Portillo Wightman et al. (2023); Quevedo et al. (2024). Therefore, we select $5$ formats as candidates in the main experiments, considering the performance of our two methods.
4.3.4 How to further improve the classification performance of FlexTaF-Single?
Format WikiTQ TabFact
Llama3 DeepSeek-Coder Llama3 DeepSeek-Coder
8B 70B 6.7B 33B 8B 70B 6.7B 33B
Markdown $78.7$ $73.8$ $47.1$ $20.3$ $64.3$ $67.1$ $25.3$ $36.0$
Dict $27.9$ $23.8$ $35.1$ $46.7$ $25.0$ $14.3$ $37.5$ $36.0$
List $3.4$ $29.6$ $9.7$ $19.4$ $20.0$ $25.0$ $0.0$ $0.0$
Pandas $2.1$ $10.0$ $28.6$ $30.1$ $0.0$ $11.1$ $22.7$ $7.7$
Database $14.3$ $15.6$ $50.6$ $29.7$ $18.5$ $26.7$ $31.3$ $29.6$
Overall $69.6$ $82.2$ $68.5$ $73.1$ $79.6$ $88.8$ $74.3$ $80.6$
Table 5: The overall classification accuracy of FlexTaF-Single using four LLMs on two datasets, and the classification accuracy for instances that can only be correctly solved by a single format, with the best performance marked in bold.
We analyze both the overall accuracy and the accuracy on instances that can be correctly solved by only one format, as these instances most distinctly highlight the unique features of each format. The classification results are presented in Table 5. We observe that: (i) The classification performance of predicting larger-scale LLMs is better than that of smaller-scale LLMs. This is attributed to the greater robustness of larger-scale LLMs Howe et al. (2024), which results in more consistent features in instances correctly solved with the same format. (ii) The classification performance of each format is positively correlated with the size of its training data (see the supplementary material). Therefore, future improvements in classification performance can be achieved by reducing noise in the data and increasing the scale of the training data.
4.3.5 How does the data filtering strategy affect FlexTaF-Single?
1111222233334444555559595959616161616363636365656565676767676969696971717171Maximum threshold of the number of labelsAccuracy of ClassificationAccuracy of ClassificationAccuracy of FlexTaF-Single41414141434343434545454547474747494949495151515153535353Accuracy of FlexTaF-Single
Figure 5: The accuracy of classification and FlexTaF-Single, with the change of the maximum threshold of the number of labels in the training data.
We compare various strategies by adjusting the maximum threshold of the number of labels in each training data instance. The experimental results are shown in Figure 5. We find that: (i) As the maximum label count rises, overall performance improves, peaking at $3$ labels, which demonstrates that training with $\leq 3$ labels per instance works best, because the increased amount of training data aids the model training. (ii) When the label count goes above $3$, performance drops. This shows why filtering training data is important, as instances easily resolved by most tabular formats are often simpler and fail to effectively capture the unique features of each label, thus negatively impacting model training.
4.4 Case Study
Figure 6: An instance from the WikiTQ test set using Llama3-70B with Markdown and Dict tabular formats.
To better illustrate that different instances are suitable for different tabular formats, we present an instance from WikiTQ. As illustrated in Figure 6, when utilizing the Dict format, Llama3-70B Meta (2024) generates an incorrect program that processes all table rows without excluding a special ”-” row. Conversely, when solving the instance with Markdown, the model successfully ignores the ”-” line and correctly identifies the country that won the most gold medals. Therefore, for this instance, the Markdown format proves more suitable than Dict when employing Llama3-70B. Additional instances are provided in the supplementary material.
5 Related Works
The table reasoning task aims to answer the natural language question based on provided tabular data (Dong et al., 2022; Zhang et al., 2024d; Dong & Wang, 2024). LLM-based methods have become predominant in this field due to their superior performance of table understanding, commonsense reasoning, and logical reasoning (Wei et al., 2022; Chen, 2023; Zhao et al., 2023; Chen et al., 2023; Gao et al., 2023). Some existing works focus on enhancing table reasoning ability by fine-tuning LLMs (Zhang et al., 2024b; Bian et al., 2024; Zhang et al., 2024c; Patnaik et al., 2024; Wu & Feng, 2024; Gardner et al., 2024; Li et al., 2024). Also, some works improve table reasoning performance by designing prompts (Zhao et al., 2023; Ye et al., 2023; Nahid & Rafiei, 2024a; b; Lee et al., 2024; Wang et al., 2024; Zhao et al., 2024) or aggregating diverse results (Ni et al., 2023; Liu et al., 2024a). For instance, Mix Self-Consistency (Liu et al., 2024a) proposes combining textual and symbolic reasoning results.
Previous works generally employ a fixed tabular format across varying instances, regardless of the specific LLM used, which could limit the performance. Since Sui et al. (2024) demonstrate that the performance of table understanding tasks, such as Cell Lookup and Size Detection, varies with different tabular formats, and propose self-augmented prompting. To aid LLMs in understanding tables, self-augmented prompting employs LLMs to extract critical table contents and summarize them into natural language descriptions for LLMs. Building on this, Singha et al. (2023) evaluate the performance of additional tabular formats and investigate the effects of noise operations like Shuffle Columns and Transpose Table on table understanding. They analyze that the LLM adopts different tabular formats, causing varying robustness in table understanding tasks. Similarly, Deng et al. (2024) examine the different performance of text-based and image-based formats in table reasoning tasks.
However, existing studies primarily discuss the impact of different tabular formats from the perspective of tasks, ignoring the impact of the instances and LLMs. Therefore, we claim that different instances and models suit different tabular formats by analyzing the experimental results. Based on the claim, we propose FlexTaF to improve the table reasoning performance by flexibly employing tabular formats.
6 Conclusion
In this paper, we explore the impact of tabular format on table reasoning performance from two aspects. (i) We claim from the perspective of experimental results that different instances and models have different most suitable tabular formats. (ii) We propose our methods. FlexTaF-Single use the classifier to predict the most suitable tabular format according to the instance and the LLM. FlexTaF-Vote obtains the results from multiple formats and employs the voting mechanism to get the final result. We build experiments on WikiTableQuestions and TabFact datasets. Compared with the best performance of using a fixed tabular format with greedy decoding and self-consistency, FlexTaF-Single and FlexTaF-Vote increase by $2.3\%$ and $4.8\%$ on average respectively, demonstrating the effectiveness of our methods.
Appendix A Introduction to Five Tabular Formats
In this section, we introduce the five tabular formats used in FlexTaF: Markdown, Dict, List, Pandas, and Database. The Markdown format refers to the representing tables in the Markdown language. The Dict format employs List[Dict: [str, Any]] to index the table, in which each row is stored in a dictionary, and the column name of each column is indexed as the key of the dictionary. The List format adopts List[List[Any]] to index the table, in which each row, including the header, is stored in the list, and each column needs to be indexed with the serial number of the column. The Pandas format is a Python code snippet that uses the Pandas DataFrame API to define the table, which points out each line of the table and the header. Database format refers to the format of representing a table as a database, describing the column name with a CREATE statement, and listing specific values.
Appendix B Prompts with Each Tabular Format
We present the prompts used in experiments in this section.
B.1 Prompts for Main Experiments
The prompt with the Markdown format for WikiTQ.
Please answer the question with the given table, present the final result in the format ”…, so the answer is: (answer)”:
Please note that utilize the format, do not include periods. Here are some instances you may refer to:
table:
$\mid Aircraft\mid Description\mid MaxGrossWeight\mid Totaldiskarea\mid MaxdiskLoading\mid$
::::$\mid$:--$\mid$:--$\mid$:--$\mid$:--$\mid$:--$\mid$ : - - - ∣ : - - - ∣ : - - - ∣ : - - - ∣ : - - - ∣
$\mid RobinsonR-22\mid Lightutilityhelicopter\mid 1,370lb(635kg)\mid...\mid$
$\mid Bell206B3JetRanger\mid Turboshaftutilityhelicopter\mid 3,200lb(1,451kg)\mid...\mid$
$\mid CH-47DChinook\mid Tandemrotorhelicopter\mid 50,000lb(22,680kg)\mid...\mid$
utterance:
What is the max gross weight of the Robinson R-22?
answer:
To find out what is the max gross weight of Robinson R-22, we need to look at the ”Max Gross Weight” column of the table provided. According to the table, the max gross weight of the Robinson R-22 is 1,370 lb (635 kg), so the answer is: 1,370 lb (635 kg)
table:
$\mid Player\mid No.\mid Nationality\mid Position\mid YearsinToronto\mid School/ClubTeam\mid$
:::::$\mid$:--$\mid$:--$\mid$:--$\mid$:--$\mid$:--$\mid$:--$\mid$ : - - - ∣ : - - - ∣ : - - - ∣ : - - - ∣ : - - - ∣ : - - - ∣
$\mid MarkBaker\mid 3\mid UnitedStates\mid Guard\mid 1998-99\mid OhioState\mid$
$\mid MarcusBanks\mid 3\mid UnitedStates\mid Guard\mid 2009-10\mid UNLV\mid$
$\mid LeandroBarbosa\mid 20\mid Brazil\mid Guard\mid 2010-2012\mid Tilibra/Copimax(Brazil)\mid$
utterance:
How many players were with the school or club team La Salle?
answer:
To count the number of players with the school or club team La Salle, we need to look at the ”School/Club Team” column of the table provided. According to the table, there are 1 player whose School/Club Team is La Salle, so the answer is: 1
table:
$\mid Model\mid 1991\mid 1995\mid 1996\mid 1997\mid 1998\mid 1999\mid 2000\mid 2001\mid 2002\mid 2003\mid 2004\mid$
:::::::::::$\mid$:--$\mid$:--$\mid$:--$\mid$:--$\mid$:--$\mid$:--$\mid$:--$\mid$:--$\mid$:--$\mid$:--$\mid$:--$\mid$:--$\mid$ : - - - ∣ : - - - ∣ : - - - ∣ : - - - ∣ : - - - ∣ : - - - ∣ : - - - ∣ : - - - ∣ : - - - ∣ : - - - ∣ : - - - ∣ : - - - ∣
$\mid SkodaFelicia\mid 172,000\mid 210,000\mid-\mid 288,458\mid 261,127\mid 241,256\mid 148,028\mid 44,963\mid-\mid$
$\mid SkodaOctavia\mid-\mid-\mid-\mid 47,876\mid 102,373\mid 143,251\mid 158,503\mid 164,134\mid 164,017\mid 165,635\mid 181,683\mid$
$\mid SkodaFabia\mid-\mid-\mid-\mid-\mid-\mid 823\mid 128,872\mid 250,978\mid 264,641\mid 260,988\mid 247,600\mid$
is the number on skoda fabia for 1999 more or less than 1000?
answer:
To find out if the number on the Skoda Fabia for 1999 is more or less than 1000, we need to look at the data provided for the ”Skoda Fabia” in ”1999”, which is 823, that is, the number for the Skoda Fabia in 1999 is less than 1000, so the answer is: less
table:
$\mid Place\mid Rider\mid Country\mid Team\mid Points\mid Wins\mid$
:::::$\mid$:--$\mid$:--$\mid$:--$\mid$:--$\mid$:--$\mid$:--$\mid$ : - - - ∣ : - - - ∣ : - - - ∣ : - - - ∣ : - - - ∣ : - - - ∣
$\mid 1\mid SylvainGeboers\mid Belgium\mid Suzuki\mid 3066\mid 3\mid$
$\mid 2\mid AdolfWeil\mid Germany\mid Maico\mid 2331\mid 2\mid$
$\mid 3\mid JohnBanks\mid UnitedKingdom\mid CZ\mid 971\mid 0\mid$
$\mid 4\mid MarkBlackwell\mid UnitedStates\mid Husqvarna\mid 604\mid 0\mid$
$\mid 5\mid BradLackey\mid UnitedStates\mid CZ\mid 603\mid 0\mid$
utterance:
which country had the most riders that placed in the top 20 of the 1971 trans-ama final standings?
answer:
To find out which country had the most riders in the top 20, we need to look at the ”Country” column of the table provided and count the number of times each country appears. According to the table, United States had the most riders, so the answer is: United States
table:
$<$table$>$
utterance:
$<$utterance$>$
answer:
Table 6: The prompt with the Markdown format used in the main experiments for WikiTQ.
The prompt with the Dict format for WikiTQ.
Answer the question with the given table using python code.
You should generate a function with the following signature without any other parameters. Here are some instances you may refer to:
table = [
        "Aircraft": "Robinson R-22",
        "Max Gross Weight": "1,370 lb (635 kg)",
utterance: What is the max gross weight of the Robinson R-22?
def solver(table):
    for row in table:
        if row["Aircraft"] == "Robinson R-22":
            return row["Max Gross Weight"]
table = [...]
utterance: How many players were with the school or club team La Salle?
def solver(table):
    players_la_salle = set()
    for row in table:
        if row["School/Club Team"] == "La Salle": players_la_salle.add(row["Player"])
    return len(players_la_salle)
table = [...]
utterance: is the number on skoda fabia for 1999 more or less than 1000?
def solver(table):
    for row in table:
        if row["Model"] == "Skoda Fabia": num_1999 = row["1999"].replace(",", "")
            if int(num_1999) > 1000: return "more"
            else: return "less"
    return "less"
table = [...]
utterance: which country had the most riders that placed in the top 20 of the 1971
trans-ama final standings?
def solver(table):
    country_counts = {}
    for row in table:
        country = row["Country"]
        if country in country_counts: country_counts[country] += 1
        else: country_counts[country] = 1
    max_riders = max(country_counts.values())
    countries_with_max_riders = [country for country, count in country_counts.items()
    if count == max_riders]
    return countries_with_max_riders[0]
Based on the above demonstrations, answer the following utterance with the following table using Python code.
table = $<$table$>$
utterance: $<$utterance$>$
def solver(table):
# Your code here
Table 7: The prompt with the Dict format used in the main experiments for WikiTQ. Due to the limited length of the paper, we do not list the contents of all the tables in the demonstrations, which are the same tables in Table 6.
The prompt with the List format for WikiTQ.
Answer the question with the given table using python code.
You should generate a function with the following signature without any other parameters. Here are some instances you may refer to:
table = [
        "Aircraft",
        "Robinson R-22",
utterance: What is the max gross weight of the Robinson R-22?
def solver(table):
    for row in table[1:]:
        if row[0] == "Robinson R-22": return row[2]
table = [...]
utterance: How many players were with the school or club team La Salle?
def solver(table):
    players_la_salle = set()
    for row in table[1:]:
        if row[5] == "La Salle": players_la_salle.add(row[0])
    return len(players_la_salle)
table = [...]
utterance: is the number on skoda fabia for 1999 more or less than 1000?
def solver(table):
    for row in table[1:]:
        if row[0] == ’Skoda Fabia’:
            if row[6] == "-" or int(row[6].replace(",", "")) < 1000: return "less"
            else:mreturn "more"
table = [...]
utterance: which country had the most riders that placed in the top 20 of the 1971
trans-ama final standings?
def solver(table):
    country_counts = {}
    for row in table[1:]:
        country = row[2]
        if country in country_counts: country_counts[country] += 1
        else: country_counts[country] = 1
    max_riders = max(country_counts.values())
    countries_max = [c for c, count in country_counts.items() if count == max_riders]
    return countries_max[0]
Based on the above demonstrations, answer the following utterance with the following table using Python code.
table = $<$table$>$
utterance: $<$utterance$>$
def solver(table):
# Your code here
Table 8: The prompt with the List format used in the main experiments for WikiTQ. Due to the limited length of the paper, we do not list the contents of all the tables in the demonstrations, which are the same tables in Table 6.
The prompt with the Pandas format for WikiTQ.
Answer the question with the given table using python code.
You should generate a function with the following signature without any other parameters. Here are some instances you may refer to:
table = pd.DataFrame([
        "Robinson R-22",
], columns = [
    "Aircraft",
utterance: What is the max gross weight of the Robinson R-22?
def solver(table):
    import pandas as pd
    max_gross_weight_r22 = table[table["Aircraft"] == "Robinson R-22"]
    ["Max Gross Weight"].iloc[0]
    return max_gross_weight_r22
table = [...]
utterance: How many players were with the school or club team La Salle?
def solver(table):
    import pandas as pd
    la_salle_count = table[table["School/Club Team"] == "La Salle"].shape[0]
    return la_salle_count
table = [...]
utterance: is the number on skoda fabia for 1999 more or less than 1000?
def solver(table):
    import pandas as pd
    fabia_row = table[table[’Model’] == ’Skoda Fabia’]
    fabia_1999_sales = fabia_row[’1999’].values[0]
    if fabia_1999_sales == ’-’ or int(fabia_1999_sales.replace(",", "")) < 1000:
        return ’less’
    else:
        return ’more’
table = [...]
utterance: which country had the most riders that placed in the top 20 of the 1971
trans-ama final standings?
def solver(table):
    import pandas as pd
    most_riders_country = table[’Country’].value_counts().idxmax()
    return most_riders_country
Based on the above demonstrations, answer the following utterance with the following table using Python code.
table = $<$table$>$
utterance: $<$utterance$>$
def solver(table):
# Your code here
Table 9: The prompt with the Pandas format used in the main experiments for WikiTQ. Due to the limited length of the paper, we do not list the contents of all the tables in the demonstrations, which are the same tables in Table 6.
The prompt with the Database format for WikiTQ.
Please complete the sql below to solve the question with the given database.
Here are some instances you may refer to:
database:
CREATE TABLE information (
year int ,
division int ,
Columns and instances in each column :
year: 1998, 1999, 2000, 2001, 2002, 2003, 2004, 2005, 2006, 2007, 2008, ... ;
utterance:
when was the first time the kansas city brass qualified for the playoffs?
sql:
SELECT year FROM information WHERE playoffs != ’Did not qualify’ ORDER
BY year ASC LIMIT 1;
database:
CREATE TABLE information (...);
utterance:
what was the next episode after \"do-si-do?\"
sql:
SELECT episode FROM information WHERE num = (SELECT num FROM information
WHERE episode = ’Do-Si-Do’) + 1;
database:
CREATE TABLE information (...);
utterance:
which dino album yielded the most songs on the billboard hot 100?
sql:
SELECT album FROM information WHERE chart = ’Billboard Hot 100’ GROUP BY album
ORDER BY COUNT(*) DESC LIMIT 1;
database:
CREATE TABLE information (...);
utterance:
when was the last year team penske finished first?
sql:
SELECT MAX(year) FROM information WHERE team = ’Team Penske’ AND finish = 1;
Based on the above demonstrations, answer the following utterance with the following database using SQL.
database:
$<$table$>$
utterance:
$<$utterance$>$
sql:
SELECT
Table 10: The prompt with the Database format used in the main experiments for WikiTQ. Due to the limited length of the paper, we do not list the contents of all the tables in the demonstrations.
The prompt with the Markdown format on TabFact.
Verify the consistency between the table and the utterance.
Please present the final result in the format ”…, so the answer is: (answer)” and the ”(answer)” is ”True” or ”False”.
Please note that utilize the format, do not include periods.
Here are some demonstrations you may refer to:
table:
$\mid tournament\mid wins\mid top-5\mid top-10\mid top-25\mid events\mid cutsmade\mid$
::::::$\mid$:--$\mid$:--$\mid$:--$\mid$:--$\mid$:--$\mid$:--$\mid$:--$\mid$ : - - - ∣ : - - - ∣ : - - - ∣ : - - - ∣ : - - - ∣ : - - - ∣ : - - - ∣
$\mid masterstournament\mid 0\mid 1\mid 2\mid 4\mid 4\mid 4\mid$
$\mid usopen\mid 0\mid 2\mid 3\mid 4\mid 6\mid 5\mid$
$\mid theopenchampionship\mid 1\mid 2\mid 2\mid 2\mid 3\mid 3\mid$
$\mid pgachampionship\mid 0\mid 0\mid 1\mid 2\mid 5\mid 4\mid$
$\mid totals\mid 1\mid 5\mid 8\mid 12\mid 18\mid 16\mid$
utterance:
tony lema be in the top 5 for the master tournament , the us open , and the open championship
answer:
To verify whether tony lema be in the top 5 for the master tournament , the us open , and the open championship, we need to look at the ”top - 5” column of the table provided. According to the table, the ”top - 5” column of the ”masters tournament”, ”us open”, and ”the open championship” are all more than zero, so the answer is: True
table:
$\mid year\mid competition\mid venue\mid position\mid event\mid$
::::$\mid$:--$\mid$:--$\mid$:--$\mid$:--$\mid$:--$\mid$ : - - - ∣ : - - - ∣ : - - - ∣ : - - - ∣ : - - - ∣
$\mid 2006\mid worldcrosscountrychampionships\mid fukuoka,japan\mid 10th\mid individualjuniorrace\mid$
$\mid 2006\mid worldcrosscountrychampionships\mid fukuoka,japan\mid 3rd\mid teamjuniorrace\mid$
$\mid 2006\mid africanchampionshipsinathletics\mid bambous,mauritius\mid 5th\mid 10000m\mid$
$\mid 2006\mid worldroadrunningchampionships\mid debrecen,hungary\mid 7th\mid individual20km\mid$
$\mid 2006\mid worldroadrunningchampionships\mid debrecen,hungary\mid 3rd\mid team20km\mid$
$\mid 2007\mid worldcrosscountrychampionships\mid mombasa,kenya\mid 7th\mid individual\mid$
$\mid 2007\mid all-africagames\mid algiers,algeria\mid 2nd\mid 10000m\mid$
$\mid 2007\mid worldchampionshipsinathletics\mid osaka,japan\mid 13th\mid 10000m\mid$
$\mid 2009\mid worldcrosscountrychampionships\mid amman,jordan\mid 17th\mid individual\mid$
$\mid 2013\mid worldchampionships\mid moscow,russia\mid 3rd\mid marathon\mid$
utterance:
japan and hungary host the competition 3 time each
answer:
To verify whether japan and hungary both host the competition 3 time, we need to look at the ”venue” column of the table provided. According to the table, ”japan” hosts the competition 3 times, but ”hungary” hosts the competition 2 times, so the answer is: False
Based on the above demonstrations, Verify the consistency between the following table and utterance.
table:
$<$table$>$
utterance:
$<$utterance$>$
answer:
Table 11: The prompt with the Markdown format used in the main experiments for TabFact.
The prompt with the Dict format for TabFact.
Verify the consistency between the table and the utterance with ”True” or ”False” using python code.
You should generate a function with the following signature without any other parameters:
Here are some demonstrations you may refer to:
table = [
        "tournament": "masters tournament",
        "wins": "0",
        "top - 5": "1",
        "top - 10": "2",
        "top - 25": "4"
        "tournament": "us open",
        "wins": "0",
        "top - 5": "2",
        "top - 10": "3",
        "top - 25": "4"
utterance: tony lema be in the top 5 for the master tournament , the us open ,
and the open championship
def solver(table):
    top_5_tournament = [row["tournament"] for row in table if int(row["top - 5"]) > 0]
    if "masters tournament" not in top_5_tournament:
        return False
    if "us open" not in top_5_tournament:
        return False
    if "the open championship" not in top_5_tournament:
        return False
    return True
table = [
        "year": "2006",
        "competition": "world cross country championships",
        "venue": "fukuoka , japan",
        "position": "10th",
        "event": "individual junior race"
utterance: japan and hungary host the competition 3 time each
def solver(table):
    japan_host_time = 0
    hungary_host_time = 0
    for row in table:
        if "japan" in row["venue"]:
            japan_host_time += 1
        elif "hungary" in row["venue"]:
            hungary_host_time += 1
    return (japan_host_time == 3 and hungary_host_time == 3)
Based on the above demonstrations, Verify the consistency between the following table and utterance.
table = $<$table$>$
utterance: $<$utterance$>$
def solver(table):
# Your code here
Table 12: The prompt with the Dict format used in the main experiments for TabFact.
The prompt with the List format for TabFact.
Verify the consistency between the table and the utterance with ”True” or ”False” using python code.
You should generate a function with the following signature without any other parameters:
Here are some demonstrations you may refer to:
table = [
        "tournament",
        "wins",
        "top - 5",
        "top - 10",
        "top - 25"
        "masters tournament",
        "0",
        "1",
        "2",
        "4"
utterance: tony lema be in the top 5 for the master tournament , the us open ,
and the open championship
def solver(table):
    top_5_tournament = [row[0] for row in table[1:] if int(row[2]) > 0]
    if "masters tournament" not in top_5_tournament:
        return False
    if "us open" not in top_5_tournament:
        return False
    if "the open championship" not in top_5_tournament:
        return False
    return True
table = [
        "year",
        "competition",
        "venue",
        "position",
        "event"
utterance: japan and hungary host the competition 3 time each
def solver(table):
    japan_host_time = 0
    hungary_host_time = 0
    for row in table[1:]:
        if "japan" in row[2]:
            japan_host_time += 1
        elif "hungary" in row[2]:
            hungary_host_time += 1
    return (japan_host_time == 3 and hungary_host_time == 3)
Based on the above demonstrations, answer the following utterance with the following table using Python code.
table = $<$table$>$
utterance: $<$utterance$>$
def solver(table):
# Your code here
Table 13: The prompt with the List format used in the main experiments for TabFact.
The prompt with the Pandas format for TabFact.
Verify the consistency between the table and the utterance with ”True” or ”False” using python code.
You should generate a function with the following signature without any other parameters:
Here are some demonstrations you may refer to, and you don’t need to answer the demonstrations:
table = pd.DataFrame([
        "masters tournament",
        "0",
        "1",
        "2",
        "4",
        "4",
        "4"
], columns = [
    "tournament",
    "wins",
    "top - 5",
    "top - 10",
    "top - 25",
    "events",
    "cuts made"
utterance: tony lema be in the top 5 for the master tournament , the us open ,
and the open championship
def solver(table):
    tournaments = ["masters tournament", "us open", "the open championship"]
    for tournament in tournaments:
        row = table[table[’tournament’] == tournament]
        if row.empty or int(row[’top - 5’].values[0]) == 0:
            return False
    return True
Based on the above demonstrations, answer the following utterance with the following table using Python code.
table = $<$table$>$
utterance: $<$utterance$>$
def solver(table):
# Your code here
Table 14: The prompt with the Pandas format used in the main experiments for TabFact.
The prompt with the Database format for TabFact.
Please complete the sql below to solve the question with the given database.
Here are some instances you may refer to:
CREATE TABLE information (
tournament text ,
wins int ,
top_5 int ,
top_10 int ,
top_25 int ,
events int ,
cuts_made int
Columns and instances in each column :
tournament: masters tournament, us open, the open championship, pga championship, totals ;
wins: 0, 0, 1, 0, 1 ;
top_5: 1, 2, 2, 0, 5 ;
top_10: 2, 3, 2, 1, 8 ;
top_25: 4, 4, 2, 2, 12 ;
utterance:
tony lema be in the top 5 for the master tournament , the us open ,
and the open championship
SQL:
SELECT CASE  WHEN (SELECT COUNT(*) FROM information WHERE tournament IN
(’masters tournament’, ’us open’, ’the open championship’) AND top_5 > 0) = 3
THEN ’True’ ELSE ’False’ END AS result;
CREATE TABLE information (
year int ,
competition text ,
venue text ,
position text ,
event text
Columns and instances in each column :
year: 2006, 2006, 2006, 2006, 2006, 2007, 2007, 2007, 2009, 2013 ;
competition: world cross country championships, world cross country championships, ... ;
venue: fukuoka , japan, fukuoka , japan, bambous , mauritius, debrecen , hungary, ... ;
position: 10th, 3rd, 5th, 7th, 3rd, 7th, 2nd, 13th, 17th, 3rd ;
event: individual junior race, team junior race, 10000 m, ... ;
utterance:
japan and hungary host the competition 3 time each
SQL:
SELECT CASE WHEN (SELECT COUNT(*) FROM information WHERE venue LIKE ’%japan%’) = 3 AND
(SELECT COUNT(*) FROM information WHERE venue LIKE ’%hungary%’) = 3 THEN ’True’
ELSE ’False’ END AS result;
Based on the above demonstrations, answer the following utterance with the following database using SQL.
database:
$<$table$>$
utterance:
$<$utterance$>$
sql:
SELECT
Table 15: The prompt with the Database format used in the main experiments for TabFact.
The prompts for WikiTQ Pasupat & Liang (2015) in main experiments with a single tabular format are shown in Table 6, Table 7, Table 8, Table 9, and Table 10. And the prompts for TabFact Chen et al. (2020) in main experiments with a single tabular format are shown in Table 11, Table 12, Table 13, Table 14, and Table 15. It can be found that only the demonstrations with the Database format are different among the prompts for WikiTQ. If we use the same demonstrations as other prompts, the SQL given in the prompt is too complicated to reduce the reasoning performance of the model.
B.2 Prompts for Experiments of Unifying Demonstrations
The prompt with the Database format using unified demonstrations.
Please complete the sql below to solve the question with the given database.
Here are some instances you may refer to:
database:
CREATE TABLE information (
aircraft text ,
Columns and instances in each column :
aircraft: Robinson R-22, Bell 206B3 JetRanger, CH-47D Chinook, ... ;
utterance:
What is the max gross weight of the Robinson R-22?
sql:
SELECT max_gross_weight FROM information WHERE aircraft = ’Robinson R-22’;
database:
CREATE TABLE information (
player text ,
utterance:
How many players were with the school or club team La Salle?
sql:
SELECT COUNT(*) FROM information WHERE school_club_team = ’La Salle’;
database:
utterance:
is the number on skoda fabia for 1999 more or less than 1000?
sql:
SELECT CASE WHEN CAST(column_1999 AS INTEGER) < 1000 THEN ’less’ ELSE ’more’
END AS result FROM information WHERE model = ’Skoda Fabia’;
database:
utterance:
which country had the most riders that placed in the top 20?
sql:
SELECT country, COUNT(rider) as rider_count FROM information WHERE place <= 20
GROUP BY country ORDER BY rider_count DESC LIMIT 1;
Based on the above demonstrations, answer the following utterance with the following database using SQL.
database:
$<$table$>$
utterance:
$<$utterance$>$
sql:
SELECT
Table 16: The prompt with the Database format with the unified demonstrations. Due to the limited length of the paper, we do not list the contents of all the tables in the demonstrations, which are the same tables in Table 6.
We present the prompt with the unified demonstrations in this subsection. From Table 6, Table 7, Table 8 and Table 9, it can be seen that the demonstrations used in the Markdown, Dict, List, and Pandas are unified, so we only change the prompt with the Database format, which is shown in Table 16.
Appendix C Tabular Format Distributions on Models
In this section, we show the specific process of proving that the tabular format is distributed differently on the model, which is discussed in §2.
Specifically, we take the number correctly represented on each model as the observed frequency $O_{i}$, calculate the average correct number of each format under different models as the expected frequency $E_{i}$, and calculate the Chi-square statistics as Equation 3.
$\mathcal{X}^{2}=\sum\frac{(O_{i}-E_{i})^{2}}{E_{i}}$ (3)
The corresponding degree of freedom $dof$ is $dof=(N_{m}-1)*(N_{r}-1)$, where $N_{m}$ is the number of models, and $N_{r}$ is the number of tabular formats. According to the calculated Chi-square $\mathcal{X}^{2}$ and the corresponding degree of freedom $dof$, we find the corresponding P-value from the square distribution table, which is less than $0.05$, proving that there is a distinct difference in the distribution of the correct tabular formats on different models.
WikiTQ TabFact
batch size $128$ $128$
epoch $200$ $400$
learning rate $51e-$ 5 $65e-$ 6
max tokens $512$ $512$
training device 2×2$\times$2 × NVIDIA A100 40G GPU 2×2$\times$2 × NVIDIA A100 40G GPU
[…]
Table 17: Classification training details in FlexTaF-Single on WikiTQ and TabFact.
WikiTQ TabFact
Llama3 DeepSeek-Coder Llama3 DeepSeek-Coder
8B 70B 6.7B 33B 8B 70B 6.7B 33B
Data Size $7,247$ $5,649$ $7,848$ $5,944$ $48,073$ $26,050$ $45,194$ $29,619$
Table 18: The training data size employed for training the classifier in FlexTaF-Single.
Appendix D Details of Training the Classifier in FlexTaF-Single
Figure 7: The proportion of each tabular format in the training data.
In this section, we introduce the details of training the classifier in FlexTaF-Single. Our model is implemented with PyTorch Paszke et al. (2019) and transformers Wolf et al. (2020). We present training details in Table 17, and we summarize the training data size employed for training the classifier to predict each LLM across various datasets in Table 18. We count the proportion of each tabular format in the training data for different LLMs and datasets, as shown in Figure 7. Specifically, to compare the proportions of different tabular formats in the training data, we calculate the proportion of each format to all labels of all instances in the training data.
Appendix E Overlap between Tabular Formats
Figure 8: The overlap among instances that are correctly solved with each tabular format using four LLMs on two datasets. The value represents the proportion that can be solved by the tabular format corresponding to the vertical axis in the instances that are solved by the format corresponding to the horizontal axis.
In this section, we analyze the overlap between instances correctly solved by different tabular formats, which is shown in Figure 8. The overlap between instances solved by different tabular formats is all $\leq 100\%$, which proves that different instances are suitable for different tabular formats. We can find that: (i) The overlap caused by using DeepSeek-Coder is greater than that caused by Llama3 because DeepSeek-Coder is better at code formats. The difference between code formats such as Dict and List is smaller than that between code formats and natural language formats, such as Markdown. (ii) The overlap caused by employing large-scale LLMs is more serious than that of small-scale LLMs because the large-scale LLMs have higher table reasoning performance, leading to more instances correctly solved by tabular formats. (iii) The overlap on TabFact is greater than that on WikiTQ because the questions of TabFact are simpler and easier to solve by more tabular formats.
Appendix F Comparison FlexTaF with Oracle
Tabular Format WikiTQ TabFact
Llama3 DeepSeek-Coder Llama3 DeepSeek-Coder
8B 70B 6.7B 33B 8B 70B 6.7B 33B
FlexTaF-Single $50.5$ $69.1$ $46.3$ $54.5$ $77.0$ $87.1$ $70.9$ $78.3$
FlexTaF-Vote $55.7$ $69.9$ $51.4$ $60.9$ $80.3$ $88.5$ $77.9$ $84.4$
Oracle $71.8$ $83.5$ $67.7$ $74.6$ $96.9$ $98.1$ $95.4$ $97.1$
Table 19: The performance of FlexTaF-Single, FlexTaF-Vote, and Oracle, which denotes the result that each instance uses the tabular format that can get the correct answer.
In this section, we compare the performance of FlexTaF with that of Oracle, as shown in Table 19. We observe that: (i) The excellent oracle performance shows that there exist differences between different formats, which further suggests that different instances have their suitable tabular formats. (ii) The performance of FlexTaF-Single is limited by classification and has a gap with the oracle performance, since the Pre-trained Language Model (PLM) cannot predict the behavior of LLMs well Qin et al. (2024); Liu et al. (2024b); Bowman (2023). However, we do not conduct experiments to fine-tune LLMs for classification limited by computing resources. (iii) The performance of FlexTaF-Vote also has a gap with the oracle performance, because the voting mechanism we adopt does not make full use of the rich semantic information in the LLM solutions, causing the limited improvement Ni et al. (2023).
Appendix G Cases Study
Figure 9: Two instances of WikiTQ test set using Llama3-70B and DeepSeek-Coder-33B with Markdown and Dict tabular format.
To clearly show the impact of the table, we show more instances in WikiTQ in this section, as shown in Figure 9. We observe that: (i) For the first instance with the question ”what is the highest points scored by an opponent?”, DeepSeek-Coder-33B correctly solves the instance with the Dict format, while Llama3-70B is suitable for Markdown and Pandas formats. (ii) When we use the same model Llama3-70B, the last instance with the question ”in cycle 4, how many contestants are older than 20?” is solved correctly with Dict, which is different from the previous instance. The cases claim that different instances and LLMs have different most suitable tabular formats.